# A Wavefield Correlation Approach to Improve Sound Speed Estimation in Ultrasound Autofocusing

### v1

@@ II-B2 Total Variation Regularization
Since pulse-echo sound speed estimation is ill-posed, minimizing the CMPE by itself is not sufficient to obtain a reasonable local sound speed estimate. To regularize the objective function for the optimization problem, a weighted form of anisotropic total variation is used
$\mathcal{R}_{TV}=\sum_{i=1}^{M}\sum_{j=1}^{N}\alpha_{x}|D_{x}(i,j)|+\alpha_{z}|D_{z}(i,j)|,$ (12)
@@
$D_{x}(i,j)=c(x_{i},z_{j})-c(x_{i-1},z_{j}),$ (13)
is the first difference of the discrete sound speed map, $c$, in the lateral dimension $x$, containing $M$ samples;
$D_{z}(i,j)=c(x_{i},z_{j})-c(x_{i},z_{j-1}),$ (14)
is the first difference of the sound speed map in the axial dimension, $z$, containing $N$ samples; $\alpha_{x}$ is the relative lateral difference weighting; and $\alpha_{z}$ is the relative axial difference weighting. By enforcing smoothness in the sound speed map using total variation regularization, the resulting sound speed map is relatively more homogeneous while still retaining boundary edges, which is more physically reasonable. Additionally, the image visually appears more reasonable without major distortions.
@@ III Methods
III-A Optimizer Configuration
[…]
For the split step angular spectrum propagation used during WFC beamforming, the simulation grid has a spacing of one-half the element pitch laterally and $\lambda/2$ axially, with a Butterworth anti-aliasing filter as in [8] applied laterally during the propagation process to prevent wraparound artifacts. For consistency and memory constraints, the center 128 elements were used for beamforming all data, except for the rat acquisitions described in the following subsection, which were beamformed using 160 elements to encompass a larger lateral span due to the shallow imaging depth.
III-B Experiments
The optimization pipeline above was compared to using DAS with time delays calculated using a straight-ray travel path assumption through the medium as in [32], with otherwise the same initialization values and optimizer setup as described above. An f/# of 1 was used to calculate the common midpoint phase error criterion when using DAS, and 48 points per ray were used for calculating the time delays. Performance between using the two beamformers for optimization was compared for simulation, phantom, and in vivo data with reverberation clutter and sound speed heterogeneities.
III-B1 Simulation
Fullwave 2 [22, 23] 2D FDTD simulations of a full synthetic aperture transmit sequence using a linear array were run with the transducer and transmit parameters included in Table I. Two simulations used segmented abdominal wall media with acoustical properties described in [40], shrunken axially to reduce the total wall thickness to 10 mm. The media below the wall then comprised of (a) a circular sound speed inclusion with a 5 mm diameter and 1570 m/s sound speed against a background of 1540 m/s, and (b) a two layer medium with sound speeds of 1420 m/s and 1555 m/s. The density was set to 1000 kg/m3 for the background region in (a) and the first layer in (b). The densities of the second layer and inclusion were calculated so the impedance would match the background and first layer, respectively. The abdominal wall tissue was assigned acoustical parameters based on [40]. All other regions were simulated with a constant 0.10 dB/MHz/cm attenuation and 0 nonlinearity. A Gaussian blur with radius 7.5 pixels was additionally applied to the acoustic parameter maps to reduce visible reflections below the abdominal wall. The simulations used 18 scatterers per resolution cell and had a grid spacing of 0.013 mm2.
@@ III-B2 Phantom
To evaluate performance on phantom data, full synthetic aperture linear array acquisitions of two types of phantoms using a Verasonics Vantage 256 scanner (Kirkland, WA) were evaluated. A chicken layer between 10-15 mm thick was placed above an ATS 549 phantom (Norfolk, VA), reused from [7] where RF channel data was collected using a full synthetic aperture sequence from a 192 element L12-3v array with a 5.95 MHz transmit center frequency and 25 MHz sampling frequency. The ground truth sound speed in the chicken layer was measured using a piston transducer driven by a pulser-receiver, with time delays measured with an oscilloscope as described in [35]. This setup contained reverberation from the chicken-phantom boundary, as well as a sizeable sound speed difference between the chicken at 1568.5 m/s and the phantom at 1450 m/s. The acquired data were demodulated and decimated by 2 prior to beamforming due to memory constraints. For WFC, the data were then remodulated prior to the simulation process.
@@
Secondly, data from a fabricated gelatin phantom originally acquired in [4] were used to further analyze beamformer performance for heterogeneous media. Wire targets in the phantom serve as point targets, and inclusions made of a gelatin-alcohol mixture created sound speed heterogeneities. An L12-5v, 50 mm array was used to acquire the data with a 6.0 MHz transmit center frequency and 25 MHz sampling frequency.
III-B3 In Vivo
@@
Thyroid and carotid artery in vivo data from three patients was originally captured in [10] using a 128-element full synthetic aperture sequence on an L12-3v transducer with a Verasonics Vantage 256 system. A transmit center frequency of 5.95 MHz was used, with a sampling frequency of 25 MHz. Due to memory constraints, this data was demodulated and decimated by 2 prior to beamforming. These acquisitions contain reverberation clutter from superficial tissue that is visible in the vessel and anechoic nodules.
@@
Additional rat liver data from [35] was used to analyze sound speed estimation performance for tissues with larger sound speed changes. Data originally were captured with a full synthetic aperture sequence using 192 elements on an L12-3v transducer with a Verasonics Vantage 256 system. The center transmit frequency was 7.81 MHz, with a 31.25 MHz sampling frequency. The global liver sound speeds were determined ex vivo in a 37°C water tank using the same procedure as the meat-layer phantom data. Four liver lobes with sound speed differences up to approximately 30 m/s were observed.
III-B4 Evaluation Criteria
@@
The sound speed estimates and resulting beamformed images were compared between the two beamformers. For simulation and phantom data with a known ground truth sound speed, the mean (bias) and standard deviation of the error is evaluated, as well as the mean absolute error (MAE). For data containing point targets or similar small, strong reflectors, the full width at half maximum (FWHM) was used to evaluate resolution. Additionally, for the in vivo data with anechoic regions, the generalized contrast to noise ratio (gCNR) [27] was evaluated to analyze the improvement in contrast and visibility.
@@ IV-A Simulation
The local sound speed estimates for the inclusion (a) and layered media (b) under an abdominal wall are shown in Figure 2, with errors for the regions underneath the abdominal wall quantified in Table II. Most noticeably, the different sound speed regions are visually more distinguished with WFC compared to the straight-ray model with DAS. The reduction in bias with WFC is especially large for the circular sound speed inclusion in simulation (a) and the first, low sound speed layer in simulation (b). Although the bottom sound speed layer in (b) has a 10 m/s higher bias for the WFC estimation, the sound speed variation is lower compared to DAS, which has a higher error standard deviation. High sound speed corners, somewhat reminiscent of shallow-depth artifacts sometimes seen in tomographic sound speed estimators [34], are visible at the top of the WFC estimates for both simulations. This corner artifact appears to focus energy inwards, away from the lateral edges of the WFC beamformed image.
Figure 2: Sound speed and final aberration-corrected image comparisons for abdominal wall simulations with a sound speed inclusion (a) and two sound speed layers (b), with the ground truth simulated sound speed shown on the leftmost panels. The sound speed estimates using WFC show clearer differentiation of different sound speed regions, with the abdominal wall, layers, and circular inclusion more accurately represented in the estimates compared to the DAS predictions.
@@ IV-B Phantom
The sound speed estimates for the chicken overlying the ATS 549 phantom are shown in Figure 3, with errors listed in Table III. The measured sound speeds for the meat and phantom regions are visualized on the leftmost plot in Figure 3, with the depth around the boundary excluded (shown in white) due to the exact boundary location being unknown. Similarly to the abdominal wall simulations, the different sound speed regions are more visibly distinguishable in the WFC estimate compared to DAS, where the chicken region is more difficult to discern from the phantom region. Quantitatively, the bias is much higher in the chicken region for DAS than WFC. The standard deviation in the WFC sound speed estimation error, however, is higher than DAS due to the estimated sound speed decreasing in value towards the meat-phantom boundary. In the phantom region, which contains reverberation clutter from the chicken layer and phantom boundary, the DAS estimate has more than 10 m/s of bias, similar to previous sound speed estimation literature [17]. Conversely, the WFC estimate has lower bias in this region, although the standard deviation is slightly higher due to the sound speed dips visible around 15 to 20 mm in depth.
Figure 3: Sound speed estimations for an acquisition with chicken over a constant sound speed phantom. The ground truth sound speeds based on calibration measurements or phantom specifications are shown on the left. A small region is excluded about the boundary between the chicken and phantom (shown in white). The chicken region appears clearly in the WFC sound speed estimation, with much lower bias compared to the DAS estimate, and the phantom region has lower error using WFC compared to DAS.
@@
The estimates for the fabricated alcohol-gelatin phantom are shown in Figure 4. Greater differentiation between the background and inclusion sound speeds are apparent in the WFC sound speed estimation, which has approximately a 25 m/s higher difference in sound speed than the DAS estimate. The inclusion in the sound speed map also appears more circular in the WFC estimate compared to the DAS estimate. The elliptic eccentricity (ratio of the major axis to the minor axis) of the inclusion region was determined based on the region retained after thresholding the sound speed at 6 standard deviations above the estimated background sound speed. The central lesion eccentricity is 0.68 for DAS and 0.27 for WFC. Since a higher eccentricity represents a more oblong elliptical shape, the WFC inclusion is more circular in shape and more closely matches the expected inclusion shape. We have previously also observed the oblong shape of inclusions with the straight-ray model in our prior implementations [32]. Although the ground truth sound speeds for the regions were not measured, the effect of the difference in sound speed estimation can be observed in the final beamformed images of DAS and WFC, indicating, most likely, improved accuracy in the WFC estimation. Compared to using an initial homogeneous sound speed, the point targets when using a local sound speed estimate when beamforming have visible resolution improvements. However, the point targets on average have 0.24 mm lower lateral FWHM when using the WFC estimate compared to using the DAS estimate. Additionally, the point target brightness is relatively more bright compared to the background using the WFC sound speed estimate compared to the DAS estimate. The background speckle brightness, excluding the point targets, is approximately -40.9 dB for the initial constant sound speed beamformed image, -45.2 dB for the DAS-estimate beamformed image, and -47.0 dB for the WFC-estimate beamformed image. The point target resolution and intensity improvements indicate that the WFC-estimated sound speed map has relatively higher accuracy compared to the DAS-estimated sound speed map.
Figure 4: Sound speed estimation and aberration correction comparisons for a fabricated alcohol-gelatin phantom. The difference between the inclusion and background sound speeds is around 25 m/s higher in the WFC prediction compared to the DAS prediction, and the background sound speed is more uniform for WFC compared to DAS. While the point target resolution improves with aberration correction from using a local sound speed map compared to using a constant sound speed, the final image using the WFC-estimated sound speed map leads to better lateral point target resolution compared to using the DAS-estimated sound speed.
@@ IV-C In Vivo
Representative sound speed estimates and beamformed image comparisons for thyroid acquisitions from two subjects at different acquisition locations are shown in Figure 5. Although the ground truth sound speeds are unknown, the images using local sound speed maps show resolution and contrast improvements from aberration correction compared to beamforming with a constant sound speed that was used to initialize the optimization process. Contrast improvements are detailed in Table IV, with regions of interest indicated by arrows on the corresponding figures. The gCNR comparisons show that WFC results in the best contrast in hypoechoic and anechoic regions. Additionally, using the estimated sound speed maps results in improved resolution in the final aberration-corrected images. In Figure 5(a), the scatterer in the center of the third inclusion has a lateral -6 dB width of 1.39 mm when beamforming with a constant sound speed. Using the DAS sound speed during beamforming results in a 0.72 mm -6 dB width, and using the WFC sound speed results in a 0.69 mm width. Similarly in Figure 5(b), the scatterers surrounding the structure in the green box have the best resolution for the image beamformed using the WFC sound speed estimation, with improvements in lateral resolution as much as 0.14 mm compared to using a constant sound speed and 0.03 mm compared to using the DAS sound speed prediction.
Figure 5: Sound speed estimates and aberration corrected images for thyroid acquisitions from different subjects. Regions of interest for gCNR calculations are indicated by arrows. The contrast in hypoechoic and anechoic regions improves more greatly for the WFC model to compared to the straight-ray DAS model. Some stronger scatterers exhibit improvements to resolution, especially when compared to using a constant sound speed to beamform, such as the scatterer in inclusion 3 in (a) or the scatterers forming the boundary of the structure in the green box in (b).
@@
Estimations for a rat liver acquisition with more tissue heterogeneity are shown in Figure 6. The measured global sound speed of the liver was 1562 m/s, resulting in an overall bias of -40 m/s for the DAS sound speed estimate and 11 m/s for the WFC sound speed estimate. The WFC estimate appears to distinguish different tissue regions more clearly than the DAS estimate. The abdominal wall above 5 mm appears more clearly as a different sound speed region from the rest of the liver. Additionally, the tissue region on the upper left appears as an entire region with lower sound speed in the WFC estimate, while only part of that region in the upper center of the liver has a lower sound speed in the DAS estimate. The aberration-corrected WFC image demonstrates improved image quality, especially on the left side, compared to the DAS image. The largest hypoechoic region around -10 mm laterally and 15 mm axially has a gCNR of 0.93 for the WFC image and a gCNR of 0.87 for the DAS image. The -6 dB lateral width for the stronger scatterers between -5 and -10 mm laterally and 10 and 15 mm axially decreases up to 0.18 mm in the WFC image compared to the DAS image. Even on the upper right, the strong scatterer at 6 mm laterally and 4 mm axially shows a slight lateral -6 dB resolution improvement of 0.06 mm for WFC relative to DAS.
Figure 6: Sound speed estimation and aberration correction comparisons for an in vivo rat liver acquisition. The abdominal wall appears more clearly as a different sound speed from the liver region in the WFC sound speed estimate compared to the DAS estimate, and most of the tissue structure visible in the upper left of the beamformed images appears as a distinct region with lower sound speed in the WFC estimate, compared to the DAS estimate that only shows a section centered at approximately 0 mm laterally as a lower sound speed. The WFC aberration corrected image shows improved resolution and contrast, especially on the left half.

### v2

@@ II-B2 Total Variation Regularization
Because pulse-echo sound speed estimation is ill-posed, minimizing the CMPE by itself is not sufficient to obtain a physically meaningful local sound speed estimate. To regularize the objective function in the optimization problem, a weighted anisotropic total variation norm is used
$\mathcal{R}_{\text{TV}}=\sum_{i=2}^{M}\sum_{j=2}^{N}\alpha_{x}|D_{x}(i,j)|+\alpha_{z}|D_{z}(i,j)|,$ (12)
@@
$\begin{split}D_{x}(i,j)&=c(x_{i},z_{j})-c(x_{i-1},z_{j}),\\ D_{z}(i,j)&=c(x_{i},z_{j})-c(x_{i},z_{j-1}),\end{split}$ (13)
denote the forward first-order finite difference operators along the $x$ and $z$ axes applied to the discrete sound speed map $c$ of size $M\times N$. In addition, $\alpha_{x}$ and $\alpha_{z}$ denote in (12) relative weights of the axial and lateral differences.
Total variation regularization is designed to favor piecewise-constant maps, with the rationale that the sound speed map obtained by minimizing Equation (6) will be homogeneous within a specific tissue, thus removing potential distortions while still retaining sharp boundaries at the transitions between different tissues.
@@ III Methods
III-A Simulations
2D Fullwave [25, 26] simulations of a full synthetic aperture transmit sequence using a linear array were employed to evaluate the proposed approach with the transducer and transmission parameters included in Table I. Two simulations were performed that used numerical phantoms containing segmented abdominal wall media with acoustical properties described in [44], shrunken axially to reduce the total wall thickness to 10 mm. The media below the wall then comprised of (a) a circular sound speed inclusion with a 5 mm diameter and 1570 m/s sound speed against a background of 1540 m/s, and (b) a two layer medium with sound speeds of 1420 m/s and 1555 m/s. The density was set to 1000 kg/m3 for the background region in simulation (a) and the first layer in simulation (b). The densities of the second layer and inclusion were calculated so the impedance would match the background and first layer, respectively, to avoid strong “coherent” reverberations. The abdominal wall tissue was assigned acoustical parameters based on [44]. All other regions were simulated with a constant 0.10 dB/MHz/cm attenuation and 0 nonlinearity. A Gaussian blur with radius 7.5 pixels was additionally applied to the acoustic parameter maps to reduce excessively large “diffuse” reverberations [12]. The simulations use 18 scatterers per resolution cell and use a grid spacing of 0.013 mm per dimension.
@@ III-A1 Phantom
To evaluate the model’s performance on phantom data, full synthetic aperture data were acquired from two types of phantoms using a Verasonics Vantage 256 scanner (Kirkland, WA) and a linear array. The first experiment re-used data from [7], where a chicken-breast layer between 10-15 mm thick was placed above an ATS 549 phantom (Norfolk, VA). RF channel data were collected from a 192-element L12-3v array with a 5.95 MHz transmit center frequency and a 25 MHz sampling frequency. The ground truth sound speed in the chicken layer was measured using a piston transducer driven by a pulser-receiver, with time delays measured with an oscilloscope as described in the Appendix of [38]. This setup contained reverberation from the chicken-phantom boundary, as well as a sizable sound speed difference between the chicken (1568.5 m/s) and the phantom (1450 m/s). The data were demodulated and decimated by 2 prior to beamforming due to reduce memory usage. For the proposed WFC model, the data were then remodulated prior to the ASM simulation process.
@@
In the second experiment, data from a fabricated gelatin phantom originally acquired for [4] were used to further analyze beamformer performance for heterogeneous media. Wire targets in the phantom serve as point targets, and cylindrical inclusions made of a gelatin-alcohol mixture created high sound speed regions relative to the background. An L12-5v, 50 mm array was used to acquire the data with a 6.0 MHz transmit center frequency and 25 MHz sampling frequency. The sound speeds of the inclusions and background were not available for these data.
III-A2 In Vivo
@@
In vivo data from [10] and [38] were also used to evaluate the proposed method. From [10], thyroid and carotid artery data were obtained from three patients acquired using a 128-element full synthetic aperture sequence on an L12-3v transducer and Vantage 256 under a Stanford IRB approved protocol. A transmit center frequency of 5.95 MHz was used, with a sampling frequency of 25 MHz. Due to memory constraints, the data were demodulated and decimated by 2 prior to beamforming. These acquisitions contain reverberation clutter from superficial tissue that is visible in the vessel and anechoic nodules.
@@
Additional rat liver data from [38] was used to analyze sound speed estimation performance in tissues with more complex sound speed distributions. Data were originally captured with a full synthetic aperture sequence using 192 elements on an L12-3v transducer. The center transmit frequency was 7.81 MHz, with a 31.25 MHz sampling frequency. The global liver sound speeds were determined ex vivo in a 37°C water tank using the same procedure as the meat-layer phantom data. In these data, speed of sound measurements from four of the liver lobes in a subset of these rats showed sound speed differences of up to 30 m/s relative to the global liver measurements.
III-B Model Implementation
[…]
For the split step angular spectrum propagation used during WFC beamforming, the simulation grid has a spacing of one-half the element pitch laterally and $\lambda/2$ axially, with a Butterworth anti-aliasing filter as in [8] applied laterally during the propagation process to prevent wraparound artifacts. For consistency and memory constraints, the center 128 elements were used for beamforming all data, except for the rat acquisitions (see Section III-A2), which were beamformed using 160 elements to encompass a larger lateral span due to the shallower imaging depth.
III-B1 Evaluation Criteria
The optimization pipeline above was compared to the approach introduced in [35], which uses DAS with time delays calculated using a straight-ray travel path assumption through the medium. We otherwise kept the initialization values and optimizer setup described above for both the reference DAS and proposed WFC experiments. An f/# of 1 was used to calculate the common midpoint phase error criterion when using DAS, and 48 points per ray were used for calculating the time delays. We compared the performance of ultrasound autofocusing with the WFC and DAS beamformers in the simulated, phantom, and in vivo data with reverberation noise and sound speed heterogeneities.
@@ III-B1 Evaluation Criteria
For simulation and phantom data with a known ground truth sound speed, the mean (bias), standard deviation of the error, and the mean absolute error (MAE) were computed. For data containing point targets or similar small, strong reflectors, the full width at half maximum (FWHM) is used to evaluate resolution. Additionally, for the in vivo data with anechoic regions, the generalized contrast to noise ratio (gCNR) [30] is evaluated as a surrogate measure of image quality.
@@ IV Results
In our results, we refer to our models as the WFC and DAS models. However, in these comparisons, the reader should keep in mind that the DAS model should really be viewed as DAS with a straight-ray model of wave propagation and not simply the delay-and-sum process.
@@ IV-A Simulation
The local sound speed estimates for the simulated inclusion (a) and layered media (b) under an abdominal wall are shown in Figure 2, with sound speed errors for the regions underneath the abdominal wall quantified in Table II. Most noticeably, the different sound speed regions are visually more distinguished with WFC compared to the straight-ray model with DAS. The reduction in bias with WFC is especially large for the circular sound speed inclusion in simulation (a) and the first, low sound speed layer in simulation (b). Although the bottom sound speed layer in (b) has a 10 m/s higher bias for the WFC estimation, the sound speed variation is reduced compared to DAS, which has a higher error standard deviation. High sound speed corners, reminiscent of shallow-depth or corner artifacts sometimes seen in tomographic sound speed estimators [37], are visible at the top of the WFC estimates for both simulations. This corner artifact appears to focus energy inwards, away from the lateral edges of the WFC beamformed image. In addition, the shape of the circular inclusion in (a) is better recovered by the WFC model. The eccentricity of the inclusion—identified as the region with a sound speed higher than 6 standard deviation above the background sound speed—is reduced from 0.58 with the DAS model to 0.42 with the WFC model. Because a higher eccentricity represents a more oblong elliptical shape, the WFC inclusion is more circular and matches more closely to the expected inclusion shape.
Figure 2: Sound speed and final aberration-corrected image comparisons for abdominal wall simulations with a sound speed inclusion (a) and two sound speed layers (b), with the ground truth simulated sound speed shown on the leftmost panels. The sound speed estimates in the WFC model show clearer spatial differentiation of the abdominal wall, horizontal layers, and circular inclusion compared to the DAS estimates.
@@ IV-B Phantom
The sound speed estimates for the chicken overlying the ATS 549 phantom are shown in Figure 3, with errors listed in Table III. The measured sound speeds for the meat and phantom regions are visualized on the leftmost plot in Figure 3, with the depth around the boundary excluded (shown in white) due to the exact location of the boundary being unknown. Similarly to the abdominal wall simulations, the different sound speed regions are more visibly distinguishable in the WFC estimate compared to DAS, where the chicken region is more difficult to discern from the phantom region. Quantitatively, the bias is much higher in the chicken region for DAS than WFC. The standard deviation in the WFC sound speed estimation error, however, is higher than DAS due to the estimated sound speed decreasing in value towards the meat-phantom boundary. In the phantom region, which contains reverberation clutter from the chicken layer and phantom boundary, the DAS estimate has more than 10 m/s of bias, similar to previous sound speed estimation literature [18]. Conversely, the WFC estimate has lower bias in this region, although the standard deviation is slightly higher due to the drop in sound speed visible at the edges around 15 to 20 mm in depth.
Figure 3: Sound speed estimations for an acquisition with chicken over a constant sound speed phantom. The ground truth sound speeds based on calibration measurements or phantom specifications are shown on the left. To avoid uncertainty in the boundary depth, a small region is excluded about the boundary between the chicken and phantom (shown in white). The chicken region appears clearly in the WFC sound speed estimation, with much lower bias compared to the DAS estimate, and the phantom region has lower error using WFC compared to DAS.
@@
The estimates for the fabricated alcohol-gelatin phantom are shown in Figure 4. Greater differentiation between the background and inclusion sound speeds are apparent in the WFC sound speed map, where the sound speed contrast is 25 m/s higher than with DAS. The inclusion in the sound speed map also appears more circular in the WFC estimate. The eccentricity of the central lesion is 0.68 for DAS and 0.27 for WFC. The oblong inclusion shape for the DAS estimation is consistent with our previous experiments with this phantom and the straight-ray model [35].
Although the ground truth sound speed map is unknown for this phantom, the effect of the difference in sound speed estimation between the two models can be observed in the wire targets in the final beamformed images. While both models improve the B-mode images, indicating more accurate sound speed modeling, the wire targets in the image based on the WFC model have an average lateral FWHM of 0.25 mm, while the DAS model has a lateral FWHM of 0.49 mm. Furthermore, the wire target brightness is higher relative to the background brightness in the WFC model. The difference between the maximum brightness of the point targets and the surrounding background is approximately 40.9 dB for the image beamformed with a constant sound speed, 45.2 dB for the image using the DAS model, and 47.0 dB for the image using the WFC model. Overall, the general improvement in image quality indicates that the WFC-estimated sound speed map has higher accuracy compared to the DAS-estimated sound speed map.
Figure 4: Sound speed estimation and aberration correction comparisons for a fabricated alcohol-gelatin phantom. The difference between the inclusion and background sound speeds is around 25 m/s higher in the WFC model compared to the DAS model, and the background sound speed is more uniform for WFC compared to DAS. The wire target resolution improves with aberration correction compared to a constant sound speed, with further improvements when WFC is used instead of DAS.
@@ IV-C In Vivo
Figure 5 depicts comparisons of representative sound speed estimates and beamformed images for thyroid acquisitions in two subjects at different scanning positions. Although the ground truth sound speeds are unknown, the images using local sound speed maps show resolution and contrast improvements from aberration correction compared to beamforming with an optimized constant sound speed (optimized based on the process in Section III-B). Contrast improvements are detailed in Table IV, with regions of interest indicated by arrows in Figure 5. The gCNR comparisons show that the WFC model achieves greater inclusion separability in hypoechoic and anechoic regions. Additionally, the WFC model results in improved resolution in the final aberration-corrected images. In Figure 5(a), the scatterer located in the center of the third inclusion has a lateral FWHM of 1.39 mm when beamforming with an optimized constant sound speed, 0.72 mm with DAS model, and 0.69 mm with the WFC model. Similarly, in Figure 5(b), the high amplitude scatterers located in the green box have a lateral FWHM of 0.38 mm with the WFC model, whereas lateral FWHM is 0.4 mm with the DAS model and 0.52 mm with the optimized constant sound speed, respectively.
Figure 5: Sound speed estimates and aberration corrected images for thyroid acquisitions from (a) Subject 1 and (b) Subject 2. Regions of interest for gCNR calculations are indicated by arrows. The contrast in hypoechoic and anechoic regions improves more greatly for the WFC model to compared to the straight-ray DAS model. The stronger scatterers exhibit improvements in resolution, especially when compared to beamforming with a constant sound speed, such as the scatterer in inclusion 3 in (a) or the scatterers forming the boundary of the structure in the green box in (b).
@@
Sound speed estimates for a rat liver acquisition with greater tissue heterogeneity are shown in Figure 6. The measured average ground truth sound speed in the liver was 1562 m/s, resulting in an overall bias of -40 m/s for the DAS sound speed estimate and 11 m/s for the WFC sound speed estimate. The WFC estimate appears to distinguish different tissue regions better than the DAS estimate, with the abdominal wall above 5 mm appearing distinct from the rest of the liver. Additionally, the liver lobe in the upper left of the B-mode image appears as a similarly-shaped region of lower sound speed in the WFC sound speed estimate, while only moderate anatomical similarity is obtained with the DAS estimate. The aberration-corrected WFC image demonstrates improved image quality compared to the DAS image as well, especially on the left side. The largest hypoechoic region indicated by a yellow arrow on the first B-mode image in Figure 6 has a gCNR of 0.93 for the WFC image and a gCNR of 0.87 for the DAS image. The lateral FWHM for the stronger scatterers indicated by green arrows between -5 and -10 mm laterally and 10 and 20 mm axially decreases up to 0.18 mm in the WFC image compared to the DAS image. On the upper right, the strong scatterer indicated by a green arrow at 6 mm laterally and 4 mm axially shows a reduction in lateral FWHM from 0.82 mm with DAS to 0.61 mm with WFC.
Figure 6: Sound speed estimation and aberration correction comparisons for an in vivo rat liver acquisition. The transition between the abdominal wall and the liver region appears more clearly in the WFC sound speed estimate compared to the DAS estimate. Most of the tissue structure visible in the upper left of the beamformed images appears as a similarly shaped region with lower sound speed in the WFC estimate. In contrast, the DAS estimate presents moderately similar anatomy with a lower sound speed centered at approximately 0 mm laterally. The WFC aberration corrected image shows improved resolution in locations indicated by green arrows and improved contrast in the region indicated by a yellow arrow.